find the midpoint of the line segment whith the given points
(2, 1), (6, 8)

Answer by checkley71(8403)

6-2=4/2=2 THUS THE X MID POINT IS 2+2=4 OR 6-2=4
8-1=7/2=3.5 THUS THE Y MID POINT IS 8-3.5=4.5 OR 1+3.5=4.5
THUS THE MIDPOINT IS (4,4.5)